OpenStudy (anonymous):

3x^2+4y=7 5x^2+2y=12 solution of x and y?

OpenStudy (anonymous):

lets try that again shall we \[3x^2+4y=7\] \[5x^2+2y=12\] double the second one get \[10x^2+4y=24\] and then since the y's are the same set \[10x^2-24=3x^2-7\] and solve for x

OpenStudy (anonymous):

\[7x^2=17\] \[x^2=\frac{17}{7}\] \[x=\pm\sqrt{\frac{17}{7}}\]

OpenStudy (anonymous):

3x^2+4y=7 10x^2+4y=24 7x^2=17 x=\[\pm \sqrt{17/7}\] 15x^2+20y=35 15x^2+6y=36 14y=-1 y=-1/14